# Performance analysis for high-dimensional Bell-state quantum illumination

### v1

@@ Acknowledgements.
The author thanks A. Pannu, A. S. Helmy, and H. El Gamal for early access to their manuscript.
@@
$\Delta_{D}=\sum_{\tilde{\bf N}}\sum_{m=1}^{M}\sum_{m^{\prime}=1}^{M}\frac{|\Delta{\tilde{\bf N}}|\sqrt{(\tilde{N}_{m}+1)(\tilde{N}_{m^{\prime}}+1)}}{[M(N_{B}+1)]^{2}}\,{}_{R}\langle{\bf e}_{m}+{\tilde{\bf N}}|\hat{\rho}^{(1)}_{RI}|{\bf e}_{m^{\prime}}+{\tilde{\bf N}}\rangle_{R},$ (87)
@@
for $M\gg N_{B}$, as will be the case for useful error probabilities. Finally, for $m\neq m^{\prime}$ we have
@@
for $M\gg N_{B}$, as will be the case for useful error probabilities. Putting together Eqs. (89) and (91) we have that the first-order correction to our $p_{D}$ approximation from Sec.  IV has essentially no effect.

### v2

@@ Acknowledgements.
The author thanks A. Pannu, A. S. Helmy, and H. El Gamal for early access to their manuscript. He also thanks R. Di Candia for alerting him to Ref. [15].
@@
$\Delta_{D}=\sum_{\tilde{\bf N}}\sum_{m=1}^{M}\sum_{m^{\prime}=1}^{M}\frac{|\Delta{\tilde{\bf N}}|\sqrt{(\tilde{N}_{m}+1)(\tilde{N}_{m^{\prime}}+1)}}{[M(N_{B}+1)]^{2}}\,{}_{R}\langle{\bf e}_{m}+{\tilde{\bf N}}|\hat{\rho}^{(1)}_{RI}|{\bf e}_{m^{\prime}}+{\tilde{\bf N}}\rangle_{R}.$ (87)
@@
for $M\gg N_{B}+1$, as will be the case in the quantum-advantage regime, cf. Fig. 1. Finally, for $m\neq m^{\prime}$ we have
@@
for $M\gg N_{B}+1$, as will be the case in the quantum-advantage regime, cf. Fig. 1. Putting together Eqs. (89) and (91) we have that the first-order correction to our $p_{D}$ approximation from Sec.  IV has essentially no effect.